8. Calculate annuity cash flows Aa Aa Your goal is to have $20,000 in your bank account by the end of four years. If the interest rate remains constant at 6% and you want to make annual identical deposits, how much will you need to deposit in your account at the end of each year to reach your goal? $3,657.48 O $3,200.30 O $4,571.85 O $5,486.22 If your deposits were made at the beginning of each year rather than an at the end, by how much would the amount of your deposit change if you still wanted to reach your goal by the end of four years? O $194.09 O $258.79 $219.97 O $323.4910. Implied interest rate and period Aa Aa E Consider the case of the following annuities and the need to compute either their expected rate of return or duration David needed money for some unexpected expenses, so he borrowed $5,464.40 from a friend and agreed to repay the loan in eight equal installments of $1,100 at the end of each year. The agreement is offering an implied interest rate of Davids friend, Keanu, has hired a financial planner for advice on retirement. Considering Keanus current expenses and expected future lifestyle changes, the financial planner has stated that once Keanu crosses a threshold of $4,991,331 in savings, he will have enough money for retirement. Keanu has nothing saved for his retirement yet, so he plans to start depositing $85,000 in a retirement fund at a fixed rate of 12.00% at the end of each year. It will take years for Keanu to reach his retirement goal.
Answer #1

Answer to Question 8.

Answer a.

Desired sum after 4 years = $20,000
Interest rate = 6%

Let annual deposit at the end of each year be $x

$x*1.06^3 + $x*1.06^2 + $x*1.06 + $x = $20,000
$x * (1.06^4 - 1) / 0.06 = $20,000
$x * 4.3746 = $20,000
$x = $4,571.85

So, annual deposit at the end of each year is $4,571.85

Answer b.

Desired sum after 4 years = $20,000
Interest rate = 6%

Let annual deposit at the beginning of each year be $x

$x*1.06^4 + $x*1.06^3 + $x*1.06^2 + $x*1.06 = $20,000
$x * 1.06 * (1.06^4 - 1) / 0.06 = $20,000
$x * 4.6371 = $20,000
$x = $4,313.06

So, annual deposit will decrease by $258.79 ($4,571.85 - $4,313.06)